OpenStudy (anonymous):

After his final exam, a student goes to the top of a building and throws his book into the air. The height h in feet of the book above the ground is given by: h = −16t 2 + 96t + 112. How long is the book in the air

OpenStudy (anonymous):

when it falls to the ground what will be the book's height??

OpenStudy (anonymous):

set this sucker = 0 and solve for t

OpenStudy (anonymous):

ok let me try it can someone else do it as well

OpenStudy (anonymous):

hmm didnt see that - so obvious!

OpenStudy (anonymous):

sure. divide by -16 and complete the square

OpenStudy (anonymous):

-16t^2 + 96t + 112=0 t^2 - 6t - 7=0 t^2 - 7t + t -7=0 t(t-7) + 1(t-7)=0 t=7 (neclect -1)

OpenStudy (anonymous):

that what i got .

OpenStudy (anonymous):

ok i got it too thanks, i may be back to check some formulas